Let X be a random variable . P(X=-2)=P(X=-1),P(X=2)=P(X=1),P(X=0)=P(X>0)=P(X<0) . obtain the probability mass function of X.


1
Expert's answer
2021-05-27T13:14:27-0400

Probability mass function:

"P(X=0)=P(X>0)=P(X<0)=1\/3"

"P(X=-2)=P(X=-1)=1\/6"

"P(X=2)=P(X=1)=1\/6"